# Apollo: Band-sequence Modeling for High-Quality Audio Restoration

### v1

@@ I Introduction
Audio restoration has gained widespread application across various scenarios, ranging from music playback to real-time communication systems. For instance, in restoring vintage music, audio restoration methods effectively rejuvenate classic music pieces eroded by time or constrained by outdated equipment [1, 2]. Moreover, these methods are found to be extensively used in speech communication, particularly in telephone or internet calls, by repairing low-quality or distorted codec audio at the receiving end, thereby delivering a clearer and more natural auditory experience [3, 4, 5, 6]. In music playback, audio restoration mitigates the degradation caused by compression, ensuring that users enjoy high-fidelity audio [3, 7, 8]. For generative models, such as those used in music generation and speech synthesis, the audio quality is crucial, and restoration methods can enhance data quality, thus significantly improving model performance [9, 10]. Robust audio restoration methods have become indispensable components of modern audio processing systems.
@@
With the rapid advancement of deep learning, NN-based methods have gradually replaced traditional signal-processing methods. Recently, GANs [12] have demonstrated substantial potential in audio super-resolution and restoration tasks [1, 13], especially in achieving high-quality restoration. In audio codecs [14, 15, 16], GANs effectively balance perceptual audio quality with distortion, offering superior restoration performance compared to traditional methods. Audio degradation typically affects the mid-to-high-frequency bands, particularly when using lossy codecs such as MP3 or AAC [17], where high-frequency information is prone to compression artifacts. An ideal generator should retain the original audio’s low-frequency components and supplement smooth and delicate mid-to-high-frequency details, thereby achieving a more realistic audio restoration effect. The Gull codec [16] has successfully demonstrated the effectiveness of GANs in the audio codec, showing significant progress in the super-resolution reconstruction of music and speech during the decoding phase of lossy codecs.
@@ II-A Overall Pipeline
Fig.1(a) presents the proposed Apollo pipeline. Apollo operates in the time-frequency domain and comprises a band-split module, a band-sequence modeling module, and a band-reconstruction module. Specifically, given compressed or distorted audio $\mathbf{S}\in\mathbb{R}^{1\times L}$, we first transfer $\mathbf{S}$ to its time-frequency domain representation $\mathbf{X}\in\mathbb{C}^{F\times T}$ using the Short-Time Fourier Transform (STFT), where $L$ denotes the length of audio, $F$ and $T$ denote the number of frequency bins and frames, respectively. Then, the band-split module maps to sub-band embeddings $\mathbf{Z}\in\mathbb{R}^{N\times T}$ using gain-shape representations $\mathbf{G}\in\mathbb{R}^{3\times M\times T}$ for each sub-band, where $N$ and $M$ denote the number of channels in sub-band embeddings and gain-shape representations, respectively. Next, the band-sequence modeling module performs joint modeling of temporal and sub-band using a stacked architecture based on Roformer [18] and temporal convolutional network (TCN) [21, 22]. Finally, the band-reconstruction module converts the output $\mathbf{Q}\in\mathbb{R}^{N\times T}$ of the band-sequence modeling module into the reconstructed complex-valued spectrogram $\mathbf{Y}\in\mathbb{C}^{F\times T}$. It uses the inverse Short-Time Fourier Transform (iSTFT) to convert $\mathbf{Y}$ to a waveform $\bar{\mathbf{S}}\in\mathbb{R}^{1\times L}$.
@@ II-C Band-sequence Modeling Module
In Apollo, we employ stacked Band-sequence modeling modules (BS modules, Fig.1(c)) to perform joint sub-band and temporal modeling with a stacking depth of $B$. Unlike BSRNN [24] and Gull [16], each BS module consists of a series of residual Roformers [18] and TCNs, which sequentially scan along the sub-band and time dimensions, and can increase the modeling capacity to improve the model performance. First, the residual Roformer is applied to the input $\mathbf{Z}$ along the frequency band dimension $K$ to obtain $\mathbf{Z}^{\prime}\in\mathbb{R}^{N\times T}$, capturing global dependencies between sub-bands while preserving the local characteristics of the frequency domain signals. Next, the TCN is applied along the time dimension $T$ on $\mathbf{Z}^{\prime}$ to generate the output $\mathbf{Q}\in\mathbb{R}^{N\times T}$. Since the $K$ sub-band features share the same feature dimension $N$, they all share a single TCN. The TCN consists of three convolutional blocks, each containing three convolutional layers. This design allows the TCN module to efficiently handle short-term dependencies and local temporal dynamics in audio signals, enhancing the model’s ability to capture and understand temporal domain features.
@@ II-D Band-reconstruction Module
II-E Training Objection
The proposed Apollo model is trained using a GAN framework to enhance the quality of audio restoration. Specifically, the discriminator network is inspired by the multi-resolution STFT discriminator, similar to the Gull codec [16]. As described in Table I, the discriminator input consists of the spectrogram’s real and imaginary parts, which are stacked into a 3D tensor along the channel dimension. To ensure energy invariance in the input, the signal is normalized to have a unit $\ell_{2}$-norm before being passed into the discriminator. The discriminator is trained using the Least Squares GAN (LSGAN) loss [25], defined as:
$L_{\text{D}}=\sum_{i=1}^{I}\mathbb{E}_{\mathbf{A}\sim p_{\text{data}}}\left[(D_{i}(\mathbf{A})-1)^{2}\right]+\sum_{i=1}^{I}\mathbb{E}_{\mathbf{Y}\sim p_{\text{G}}}\left[(D_{i}(\mathbf{Y}))^{2}\right],$ (3)
@@ II-E Training Objection
$L_{\text{rec}}=\frac{1}{W}\sum_{w=1}^{W}\frac{\left\||\text{STFT}_{w}(\mathbf{Y})|-|\text{STFT}_{w}(\mathbf{A})|\right\|_{1}}{\left\||\text{STFT}_{w}(\mathbf{T})|\right\|_{1}},$ (4)
@@
The reconstruction loss $L_{\text{rec}}$ is calculated as the multi-resolution frequency MAE over several STFT window sizes, ensuring that both short- and long-term signal characteristics are restored effectively:
$L_{\text{rec}}=\frac{1}{W}\sum_{w=1}^{W}\left\||\text{STFT}_{w}(\mathbf{Y})|-|\text{STFT}_{w}(\mathbf{A})|\right\|_{1}.$ (6)
@@ III-A Datasets
We trained and tested Apollo on the combined MUSDB18-HQ [19] and MoisesDB [20] datasets. By integrating these two datasets, we leveraged their rich diversity and comprehensive musical resources to evaluate Apollo’s restoration performance across different music genres more thoroughly. During the data preprocessing stage, inspired by music separation techniques [10, 26], we employed a Source Activity Detector (SAD) to remove silent regions from the tracks, retaining only the significant portions for training. Throughout training, we implemented real-time data augmentation by randomly mixing tracks from different songs. Specifically, we randomly selected between 1 and 8 stems from 11 available tracks and extracted 3-second clips from each selected stem. These clips were then randomly scaled in energy within a range of [-10, 10] dB relative to their original levels. All selected stem clips were summed together to generate simulated music. Subsequently, we simulated dynamic bitrate scenarios by applying MP3 codecs with bitrates of [24000, 32000, 48000, 64000, 96000, 128000] to generate the compressed music. To ensure all samples were on the same scale, we rescaled both the target audio and the encoded audio based on the maximum absolute value.
@@ III-B Hyperparameters
For the proposed Apollo model, the Short-Time Fourier Transform (STFT) window length was set to 20 ms with a hop size of 10 ms, using a Hanning window. The bandwidth for frequency band segmentation was set to 160 Hz, and the feature dimension $N$ was set to 256. The Band Sequence modeling module was stacked $B=6$ times. In the discriminator network, the STFT window sizes were configured with a multi-scale setup, including $[32,64,128,256,512,1024,2048]$. For the optimizer, both the generator and discriminator utilize the AdamW optimizer [27]. The generator’s initial learning rate was set to 0.001, with a weight decay of 0.01, while the discriminator’s initial learning rate was set to 0.0001, with the same weight decay of 0.01. The learning rate decayed by 0.98 every two epochs, and gradient clipping with a maximum norm of 5 was employed to prevent gradient explosion. Additionally, we implemented an early stopping mechanism to prevent overfitting: training was terminated if the validation loss did not decrease for 20 consecutive epochs. All models were trained on 8 Nvidia RTX 4090 GPUs.
@@ IV Results
We evaluated the restoration performance of the Stochastic-Restoration-GAN (SR-GAN) [1] and Apollo models across various bitrates and music genres on the combined test set of MUSDB18-HQ and MoisesDB (with 5000 samples for each case). The test set encompasses a wide range of music genres, including vocals, single instruments, and mixed instruments, aiming to comprehensively assess each model’s restoration capabilities.
Bitrate Impact Analysis. Fig.2 compares the performance of the Apollo model and the Stochastic-Restoration-GAN (SR-GAN) at different bitrates (ranging from 24 kHz to 128 kHz). The experimental results demonstrated that Apollo consistently outperformed SR-GAN across all bitrates, particularly in addressing issues such as frequency band voids or reduced signal bandwidth, as reflected by SI-SNR and SDR scores. Additionally, Apollo significantly improved audio restoration quality as measured by VISQOL. Project page [^2] https://cslikai.cn/Apollo/ for Apollo’s reconstructed audio given multiple MP3 bitrates.

### v2

@@ I Introduction
Audio restoration has gained widespread application across various scenarios, ranging from music playback to real-time communication systems. For instance, in restoring vintage music, audio restoration methods effectively rejuvenate classic music pieces eroded by time or constrained by outdated equipment [1, 2, 3]. Moreover, these methods are found to be extensively used in speech communication, particularly in telephone or internet calls, by repairing low-quality or distorted codec audio at the receiving end, thereby delivering a clearer and more natural auditory experience [4, 5, 6, 7, 8, 9]. In music playback, audio restoration mitigates the degradation caused by compression, ensuring that users enjoy high-fidelity audio [4, 10, 11]. For generative models, such as those used in music generation and speech synthesis, the audio quality is crucial, and restoration methods can enhance data quality, thus significantly improving model performance [12, 13, 14]. Robust audio restoration methods have become indispensable components of modern audio processing systems [15].
@@
With the rapid advancement of deep learning, NN-based methods have gradually replaced traditional signal-processing methods. Recently, GANs [17] have demonstrated substantial potential in audio generation [18], super-resolution and restoration tasks [1, 19], especially in achieving high-quality restoration. In audio codecs [20, 21, 22], GANs effectively balance perceptual audio quality with distortion, offering superior restoration performance compared to traditional methods. Audio degradation typically affects the mid-to-high-frequency bands, particularly when using lossy codecs such as MP3 or AAC [23], where high-frequency information is prone to compression artifacts. An ideal generator should retain the original audio’s low-frequency components and supplement smooth and delicate mid-to-high-frequency details, thereby achieving a more realistic audio restoration effect. The Gull codec [22] has successfully demonstrated the effectiveness of GANs in the audio codec, showing significant progress in the super-resolution reconstruction of music and speech during the decoding phase of lossy codecs.
@@ II-A Overall Pipeline
Fig.1(a) presents the proposed Apollo pipeline. Apollo operates in the time-frequency domain and comprises a band-split module, a band-sequence modeling module, and a band-reconstruction module. Specifically, given compressed or distorted audio $\mathbf{S}\in\mathbb{R}^{1\times L}$, we first transfer $\mathbf{S}$ to its time-frequency domain representation $\mathbf{X}\in\mathbb{C}^{F\times T}$ using the Short-Time Fourier Transform (STFT), where $L$ denotes the length of audio, $F$ and $T$ denote the number of frequency bins and frames, respectively. Then, the band-split module maps to sub-band embeddings $\mathbf{Z}\in\mathbb{R}^{N\times K\times T}$ using gain-shape representations $\mathbf{G}\in\mathbb{R}^{3\times M\times T}$ for each sub-band, where $N$ and $M$ denote the number of channels in sub-band embeddings and gain-shape representations, respectively. Next, the band-sequence modeling module performs joint modeling of temporal and sub-band using a stacked architecture based on Roformer [24] and temporal convolutional network (TCN) [27, 28]. Finally, the band-reconstruction module converts the output $\mathbf{Q}\in\mathbb{R}^{N\times K\times T}$ of the band-sequence modeling module into the reconstructed complex-valued spectrogram $\mathbf{Y}\in\mathbb{C}^{F\times T}$. It uses the inverse Short-Time Fourier Transform (iSTFT) to convert $\mathbf{Y}$ to a waveform $\bar{\mathbf{S}}\in\mathbb{R}^{1\times L}$.
@@ II-C Band-sequence Modeling Module
In Apollo, we employ stacked Band-sequence modeling modules (BS modules, Fig.1(c)) to perform joint sub-band and temporal modeling with a stacking depth of $B$. Unlike BSRNN [30] and Gull [22], each BS module consists of a series of residual Roformers [24] and TCNs, which sequentially scan along the sub-band and time dimensions, and can increase the modeling capacity to improve the model performance. First, the residual Roformer is applied to the input $\mathbf{Z}$ along the frequency band dimension $K$ to obtain $\mathbf{Z}^{\prime}\in\mathbb{R}^{N\times K\times T}$, capturing global dependencies between sub-bands while preserving the local characteristics of the frequency domain signals. Next, the TCN is applied along the time dimension $T$ on $\mathbf{Z}^{\prime}$ to generate the output $\mathbf{Q}\in\mathbb{R}^{N\times K\times T}$. Since the $K$ sub-band features share the same feature dimension $N$, they all share a single TCN. The TCN consists of three convolutional blocks, each containing three convolutional layers. This design allows the TCN module to efficiently handle short-term dependencies and local temporal dynamics in audio signals, enhancing the model’s ability to capture and understand temporal domain features.
@@ II-D Band-reconstruction Module
II-E Training Objective
The proposed Apollo model was trained using a GAN framework to enhance the quality of audio restoration. Specifically, the discriminator network is inspired by the multi-resolution STFT discriminator, similar to the Gull codec [22]. As described in Table I, the discriminator input consists of real and imaginary parts of the spectrogram, which are stacked into a 3D tensor along the channel dimension. To ensure energy invariance in the input, the signal was normalized to have a unit $\ell_{2}$-norm before being passed into the discriminator. The discriminator is trained using the Least Squares GAN (LSGAN) loss [31], defined as:
$L_{\text{GAN}}=\sum_{i=1}^{I}\mathbb{E}_{\mathbf{A}\sim p_{\text{data}}}\left[(D_{i}(\mathbf{A})-1)^{2}\right]+\sum_{i=1}^{I}\mathbb{E}_{\mathbf{Y}\sim p_{\text{G}}}\left[(D_{i}(\mathbf{Y}))^{2}\right],$ (3)
@@ II-E Training Objective
$L_{\text{rec}}=\frac{1}{W}\sum_{w=1}^{W}\frac{\left\||\text{STFT}_{w}(\mathbf{Y})|-|\text{STFT}_{w}(\mathbf{A})|\right\|_{1}}{\left\||\text{STFT}_{w}(\mathbf{A})|\right\|_{1}},$ (4)
@@
where $\text{sg}[\mathbf{H}_{i,j}]$ denotes $\mathbf{H}_{i,j}$ detached from the computational graph.
@@ III-A Datasets
We trained and tested Apollo on the combined MUSDB18-HQ [25] and MoisesDB [26] datasets. By integrating these two datasets, we leveraged their rich diversity and comprehensive musical resources to evaluate Apollo’s restoration performance across different music genres more thoroughly. During the data preprocessing stage, inspired by music separation techniques [13, 32], we employed a Source Activity Detector (SAD) to remove silent regions from the tracks, retaining only the significant portions for training. Throughout the training, real-time data augmentation was implemented by randomly mixing tracks from different songs. Specifically, we randomly selected between 1 and 8 stems from 11 available tracks and extracted 3-second clips from each selected stem. These clips were then randomly scaled in energy within a range of [-10, 10] dB relative to their original levels. All selected stem clips were summed to generate simulated music. Subsequently, we simulated dynamic bitrate scenarios by applying MP3 codecs [^1] https://trac.ffmpeg.org/wiki/Encode/MP3 with bitrates of [24, 32, 48, 64, 96, 128] kbit/s to generate the compressed music. To ensure that all samples were on the same scale, we rescaled both the target audio and encoded audio based on the maximum absolute value.
@@ III-B Hyperparameters
For the proposed Apollo model, the Short-Time Fourier Transform (STFT) window length was set to 20 ms with a hop size of 10 ms, using a Hanning window. The bandwidth for frequency band segmentation was set to 160 Hz, and the feature dimension $N$ was set to 256. The Band Sequence modeling module was stacked $B=6$ times. In the discriminator network, the STFT window sizes were configured with a multi-scale setup, including $[32,64,128,256,512,1024,2048]$. For the optimizer, both the generator and discriminator utilize the AdamW optimizer [33]. The generator’s initial learning rate was set to 0.001, with a weight decay of 0.01, while the discriminator’s initial learning rate was set to 0.0001, with the same weight decay of 0.01. The learning rate decayed by 0.98 every two epochs, and gradient clipping with a maximum norm of 5 was employed to prevent gradient explosion. Additionally, we implemented an early stopping mechanism to prevent overfitting: training was terminated if the validation loss did not decrease for 20 consecutive epochs. All the models were trained on eight NVIDIA RTX 4090 GPUs.
@@ IV Results
Due to the lack of openly available baselines for this task, it is not easy for us to make a fair comparison with other related works. We evaluated the restoration performance of the Stochastic-Restoration-GAN (SR-GAN) [1] and Apollo models across various bitrates and music genres on the combined test set of MUSDB18-HQ and MoisesDB (with 5000 samples for each case). The test set encompasses a wide range of music genres, including vocals, single instruments, and mixed instruments, aiming to comprehensively assess each model’s restoration capabilities.
Bitrate Impact Analysis. Fig.2 compares the performance of the Apollo model and the Stochastic-Restoration-GAN (SR-GAN) at different bitrates (ranging from 24 kbit/s to 128 kbit/s). The experimental results demonstrated that Apollo consistently outperformed SR-GAN across all bitrates, particularly in addressing issues such as frequency band voids or reduced signal bandwidth, as reflected by SI-SNR and SDR scores. Additionally, Apollo significantly improved audio restoration quality as measured by VISQOL. Project page [^3] https://cslikai.cn/Apollo/ for Apollo’s reconstructed audio given multiple MP3 bitrates.